The pressure recorded by a constant volume gas thermometer at Kelvin temperature T is 5.92 x 10⁴ Nm-2. Calculate T if the pressure at the triple point 273 K is 5.40 x 10⁴ Nm-2.
1
Expert's answer
2020-12-04T07:22:08-0500
"\frac{p_1}{T_1}=\frac{p_2}{T_2}\\\frac{54000}{273}=\frac{59200}{T_2}\\T=299\ K"